# Data sorting modes of phoswich detector array

### v1

@@ 3 Data sorting modes and detection efficiencies
The PARIS data could be arranged into six sorting modes. For each mode, it is necessary to determine the individual detection efficiency. There is one Monte Carlo simulation package for PARIS, SToGS; see the user guide [19] to install and run the program. These six possible modes are listed in Tab. 4: 1. Only LaBr3(Ce) signals and each phoswich works individually, no add-back procedure. NaI(Tl) crystals are used as veto detectors; 2. Only LaBr3(Ce) vetoed by NaI(Tl) but being added back inside one cluster, nine phoswich detectors; 3. Only LaBr3(Ce) vetoed by NaI(Tl) but being added backed totally, twenty-seven phoswich detectors of three clusters; 4. Summed spectrum from individual LaBr3 (Ce) vetoed by NaI(Tl) meaning that the full energy is deposited in LaBr3(Ce) and individual NaI(Tl); 5. Add back the signals from LaBr3 and NaI(Tl) crystals inside one phoswich considering the cross-talk signals between two crystals, the part marked with "E, F" in Fig. 2; 6. Add back the signals from LaBr3 and NaI(Tl) crystals inside one cluster, considering the cross-talk signal inside one cluster including nine LaBr3(Ce) and nine NaI(Tl) crystals. It is important to highlight that Tab. 4 delineates the different modes for sorting the data collected by the phoswich array, whereas Tab. 1 lists the energy deposition scenarios arising when a ray interacts with the array. For instance, the application of mode 2 results in an enhanced detection efficiency of the full-energy peaks corresponding to the situation B. Nevertheless, events from situations A are also present in the spectrum at the full-energy peaks. The NaI(Tl) veto was implemented by gating on the first peak illustrated in Fig. 2.
@@
Fig. 8 shows the -ray detection efficiencies for three of the above sorting modes, which were obtained through normalizing them to the detection efficiency of the clover detector. The data presented are from -delayed rays of 80Ge [9]. The numbers of emitted rays were determined by combining the clover-HPGe-measured intensities with the GEANT4-simulated efficiency (see Fig. 9). The fit function with four parameters, shown in formula 1, is taken from radware framework [20]. The Eurisys clover detector [21] is composed of four coaxial n-type Ge diodes with diameters of 50 mm and lengths of 70 mm. The distance between the end-cup of the clover detector and the source collection point is 74 mm. The added absolute full-energy peak efficiency of four crystals is used to characterize the detection efficiencies of PARIS including three clusters (27 phoswich detectors) in different modes. Compared with mode 1, the full-energy peak detection efficiency is lower in the low-energy region in mode 2 while higher in higher-energy region, above 3.5 MeV. In mode 4, i.e. summing events which deposit full energy in LaBr3(Ce) (the first strong peak in Fig. 2) and events which only deposit full/partial energy in NaI(Tl) crystal (the second peak in Fig. 2), the full-energy peak detection efficiency is higher than in mode 1 in the whole energy region.
@@
For the detection efficiency of the clover HPGe, the low-energy region was calibrated using 152Eu sources, as shown by the red curve in Fig. 9. For the high energy region, e.g., up to 10 MeV, there is no physical calibration source. We performed numerical simulations [22, 23], first reproducing the data obtained at lower energy with the 152Eu source then extending the efficiency curve to the high-energy region using same parameters. Good agreement can be found between the data collected from 152Eu source and those generated from Geant4 simulation, as shown in the inset of Fig. 9. Fig. 9 shows the efficiency as a function of -ray energy of the clover HPGe and the associated fits. The fit function with seven parameters, shown in formula 2, is taken from radware framework [20].
@@
However, due to electron-positron pair production, there are following annihilation strong SE peaks (losing one 511 keV ray) and DE peaks (losing two 511 keV rays), especially in the high-energy region above 3 MeV. We found that this phenomenon is more pronounced in HPGe detectors, possibly because of the use of clover HPGe detectors, whereas it is not significant for PARIS even in mode 1 due to the vetoing function of the outer-volume NaI(Tl) crystals.
@@
Fig. 11 presents results of imposing multiplicity condition in PARIS. It is evident that multiplicity gating produces a slight effect on escape suppression, as demonstrated by the SEs of 5354 and 5387 keV, and leads to a modest enhancement of the signal-to-noise ratio. Nevertheless, there is a loss of statistics for full-energy peaks when applying multiplicity conditions. Notably, the spectra presented in Fig. 11 are NaI(Tl) vetoed.
@@ 4 Conclusion
In conclusion, we investigated multiple data sorting modes of the PARIS phoswich detector array for high-energy -ray detection, presenting results for time resolution, energy resolution, and detection efficiencies under different modes. Compared with a 2"2"2" LaBr3(Ce) detector [1], PARIS more effectively rejects escape events via the veto function of its outer-volume NaI(Tl) crystals in both individual and add-back modes, and rejects pile-up events by selecting events with Qlong/Qshort smaller than 3. In contrast to the common approach of adding back the energies in LaBr3(Ce) and NaI(Tl) to increase the peak-to-total ratio, using NaI(Tl) as a veto shield provides a superior trade-off for applications where spectral purity is essential. Multiplicity conditions provide additional suppression of escape events or background but simultaneously reduces the statistics of full-energy peaks. Notably, the methods presented in this work refer specifically to the -decay experiment of 80g+mGa conducted with three PARIS clusters comprising 27 phoswich detectors, rather than to a general report on the PARIS array or its overall performance. The developed method opens an avenue for further studies of the performance of the phoswich detector array in different data sorting methods, which have nuclear experimental and -ray detecting satellite implications.
@@ Acknowledgements
The authors thankfully acknowledge the work of the ALTO technical staff for the excellent operation of the ISOL source. R.L. acknowledges support by the China Scholarship Council under Grant No.201804910509. C.D., A. K. and L. A. A. received funding from European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 771036 (ERC CoG MAIDEN). Use of the PARIS modular array from the PARIS Collaboration and Ge detectors from the French-UK IN2P3-STFC Gamma Loan Pool are acknowledged.

### v2

@@ 3 Data sorting modes and detection efficiencies
The PARIS data could be arranged into six sorting modes. For each mode, it is necessary to determine the individual detection efficiency. There is one Monte Carlo simulation package for PARIS, SToGS; see the user guide [23] to install and run the program. These six possible modes are listed in Tab. 4: 1-6. 1. Only LaBr3(Ce) signals and each phoswich works individually, i.e. no add-back procedure, and NaI(Tl) crystals are used as veto detectors; 2. Only LaBr3(Ce) vetoed by NaI(Tl) but being added back inside one cluster of nine phoswich detectors; 3. Only LaBr3(Ce) vetoed by NaI(Tl) but being added back totally, i.e. within twenty-seven phoswich detectors of three clusters; 4. Summed spectrum from individual LaBr3 (Ce) vetoed by NaI(Tl) meaning that the full energy is deposited in LaBr3(Ce) and individual NaI(Tl); 5. Add back the signals from LaBr3 and NaI(Tl) crystals inside one phoswich considering the cross-talk signals between two crystals, the part marked with "E, F" in Fig. 2; 6. Add back the signals from LaBr3 and NaI(Tl) crystals inside one cluster, considering the cross-talk signal inside one cluster including nine LaBr3(Ce) and nine NaI(Tl) crystals. It is important to highlight that Tab. 4 delineates the different modes for sorting the data collected by the phoswich array, whereas Tab. 1 lists the energy deposition scenarios arising when a ray interacts with the array. For instance, the application of mode 2 results in an enhanced detection efficiency of the full-energy peaks corresponding to the situation B. Nevertheless, events from situations A are also present in the spectrum at the full-energy peaks. The NaI(Tl) veto was implemented by gating on the first peak illustrated in Fig. 2.
Fig. 8 shows the -ray detection efficiencies for three of the above sorting modes, which were obtained through normalizing them to the detection efficiency of the clover detector. The data presented are from -delayed rays of 80Ge [8]. The numbers of emitted rays were determined by combining the clover-HPGe-measured intensities with the GEANT4-simulated efficiency (see Fig. 9). The fit function with four parameters, shown in formula 1, is taken from radware framework [24]. The Eurisys clover detector [25] is composed of four coaxial n-type Ge diodes with diameters of 50 mm and lengths of 70 mm. The distance between the end-cup of the clover detector and the source collection point is 74 mm. The added absolute full-energy peak efficiency of four crystals is used to characterize the detection efficiencies of PARIS including three clusters (27 phoswich detectors) in different modes. Compared with mode 1, the full-energy peak detection efficiency is lower in the low-energy region in mode 2 while higher in the higher-energy region, above 3.5 MeV. In principle, add-back sums the energies recorded simultaneously in multiple detector segments when a single ray deposits its energy in more than one segment due to Compton scattering and -escape events. In practice, however, [e.g., mode 2 (add-back within-one-cluster) used in the present work] pseudo-add-back occurs when more than one ray from a -delayed cascade hits a single cluster. This pseudo-add-back is particularly significant for low-energy rays, which are located at the bottom of the decay level scheme and therefore have more coincidences with other lines. Consequently, compared to mode 1, mode 2 exhibits a lower full-energy peak detection efficiency in the low-energy region [Fig. 8], and mode 3 similarly shows lower counts in the low-energy region, as illustrated in Fig. 10. In mode 4, i.e. summing events which deposit full energy in LaBr3(Ce) (the first strong peak in Fig. 2) and events which only deposit full/partial energy in NaI(Tl) crystal (the second peak in Fig. 2), the full-energy peak detection efficiency is higher than in mode 1 in the whole energy region.
@@
For the detection efficiency of the clover HPGe, the low-energy region was calibrated using 152Eu source, as shown by the red curve in Fig. 9. For the high-energy region, e.g., up to 10 MeV, there is no physical calibration source. We performed numerical simulations [26, 27], first reproducing the data obtained at lower energy with the 152Eu source then extending the efficiency curve to the high-energy region using same parameters. Good agreement can be found between the data collected from 152Eu source and those generated from Geant4 simulation, as shown in the inset of Fig. 9. Fig. 9 shows the efficiency as a function of -ray energy of the clover HPGe and the associated fits. The fit function with seven parameters, shown in formula 2, is taken from radware framework [24].
@@
However, due to electron-positron pair production followed by positron annihilation, there are strong SE peaks (losing one 511 keV ray) and DE peaks (losing two 511 keV rays), especially in the high-energy region above 3 MeV. We found that this phenomenon is more pronounced in HPGe detectors, possibly because of the use of clover HPGe detectors, whereas it is not significant for PARIS even in mode 1 due to the vetoing function of the outer-volume NaI(Tl) crystals.
@@
Fig. 11 presents results of imposing multiplicity conditions in PARIS. It is evident that multiplicity gating produces a slight effect on escape suppression, as demonstrated by the SEs of 5354 and 5387 keV, and leads to a modest enhancement of the signal-to-noise ratio. Nevertheless, there is a loss of statistics for full-energy peaks when applying multiplicity conditions. Notably, the spectra presented in Fig. 11 are NaI(Tl) vetoed.
@@ 4 Conclusion
In conclusion, we investigated multiple data-sorting modes of the PARIS phoswich detector array for high-energy -ray detection, presenting results for time resolution, energy resolution, and detection efficiencies under different modes. Compared with a 2"2"2" LaBr3(Ce) detector [17], PARIS more effectively rejects escape events via the veto function of its outer-volume NaI(Tl) crystals in both individual and add-back modes, and rejects pile-up events by selecting events with Qlong/Qshort smaller than 3. In contrast to the common approach of adding back the energies in LaBr3(Ce) and NaI(Tl) to increase the peak-to-total ratio, using NaI(Tl) as a veto shield provides a superior trade-off for applications where spectral purity is essential. Multiplicity conditions provide additional suppression of escape events or background but simultaneously reduces the statistics of full-energy peaks. Notably, the methods presented in this work refer specifically to the -decay experiment of 80g+mGa conducted with three PARIS clusters comprising 27 phoswich detectors, rather than to a general report on the PARIS array or its overall performance. The developed method opens an avenue for further studies of the performance of the phoswich detector array in different data-sorting methods, which have nuclear experimental and -ray detecting satellite implications.
@@ Acknowledgements
The authors thankfully acknowledge the work of the ALTO technical staff for the excellent operation of the ISOL source. R.L. acknowledges support by the China Scholarship Council under Grant No.201804910509. C.D., A.K., and L.A.A. received funding from European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 771036 (ERC CoG MAIDEN). Use of the PARIS modular array from the PARIS Collaboration and Ge detectors from the French-UK IN2P3-STFC Gamma Loan Pool is acknowledged.